### v1

Abstract
We present follow-up SCExAO/CHARIS and -band (R 70) high-contrast integral field spectroscopy and Keck/NIRC2 photometry of directly-imaged brown dwarf companion HD 33632 Ab and new radial-velocity data for the system from the SOPHIE spectrograph, complemented by Hipparcos and Gaia astrometry. These data enable more robust spectral characterization compared to lower-resolution spectra from the discovery paper and more than double the available astrometric and radial-velocity baseline. HD 33632 Ab’s spectrum is well reproduced by a field L8.5–L9.5 dwarf. Using the Exo-REM atmosphere models, we derive a best-fit temperature, surface gravity and radius of = 1250 , log(g) = 5, and = 0.97 and a solar C/O ratio. Adding the SOPHIE radial-velocity data enables far tighter constraints on the companion’s orbital properties (e.g. =o) and dynamical mass () than derived from imaging data and Gaia eDR3 astrometry data alone. HD 33632 Ab should be a prime target for multi-band imaging and spectroscopy with the James Webb Space Telescope and the Roman Space Telescope’s Coronagraphic Instrument, shedding detailed light on HD 33632 Ab’s clouds and chemistry and providing a key reference point for understanding young exoplanet atmospheres.
1 Introduction
In the past decade, large direct imaging campaigns using extreme adaptive optics (extreme AO) systems coupled with near-infrared (IR) integral field spectrographs have discovered jovian planets and brown dwarfs on solar system scales and characterized their atmospheres (e.g. Macintosh et al., 2015; Chauvin et al., 2017; Konopacky et al., 2016; Rajan et al., 2017; Cheetham et al., 2019; Currie et al., 2023b). Most of these programs choose targets in an unbiased way, selecting stars for imaging observations based on properties like age and distance (Nielsen et al., 2019; Vigan et al., 2021; Currie et al., 2023a). Unfortunately, despite surveying hundreds of stars, such “blind” surveys have resulted in a low yield of new substellar companion discoveries: e.g. two new companions out of 300 targeted with the Gemini Planet Imager Exoplanet Survey.
| UT Date | Instrument | Seeingc (″) | Filter | ()a | PA (o) | SNRd (HD 33632 Ab) | ||
| 20211018 | SCExAO/CHARISa | 0.5–0.7 | 1.47–1.79 | 60.48 | 27 | 19.28 | 27.60 | |
| 20211018 | SCExAO/CHARISa | 0.5–0.7 | 2.01–2.36 | 60.48 | 25 | 14.86 | 29.16 | |
| 20240223 | Keck/NIRC2a | 0.2-0.5 | 3.78 | 30 | 50 | 26.86 | 29.88 |
For CHARIS data, the column refers to the wavelength range. For NIRC2 imaging data, it refers to the central wavelength.
Recent studies have demonstrated the advantage of a different approach to direct imaging discovery and characterization: targeting stars showing dynamical evidence for a substellar companion from precision calibrated astrometry from the Gaia and Hipparcos missions. This strategy has led to the discovery of planets orbiting HIP 99770 and AF Lep, a planet or brown dwarf around HIP 39017, and numerous other brown dwarf companions on 10–30 au orbits (Currie et al., 2023b; De Rosa et al., 2023; Mesa et al., 2023; Tobin et al., 2024; Franson et al., 2023; Li et al., 2023; Kuzuhara et al., 2022; Swimmer et al., 2022). Discovery yields from these surveys are up to 5 times higher than those from unbiased surveys (El Morsy et al. 2024 submitted).
A focus on direct imaging these accelerating stars also enables more in-depth characterization studies. Direct imaging data by itself cannot constrain the mass of a planet or brown dwarf. Relative astrometry for most planets and brown dwarfs imaged cover a small fraction of their orbits, resulting in characteristically poor constraints on the companions’ semimajor axes, inclinations, eccentricities, and other parameters (e.g. see Bowler et al., 2020). However, jointly analyzing the relative astrometry of a companion from imaging with the absolute astrometry of the star from Gaia and Hipparcos can directly yield dynamical masses and precise orbital parameters (e.g. Brandt, 2021; Brandt et al., 2021a).
HD 33632 Ab is a 20 au-separation brown dwarf companion orbiting a Sun-like star and one of the first examples of a joint direct imaging and astrometric discovery of a substellar companion. The companion was discovered by Currie et al. (2020a) from direct imaging and spectroscopy data obtained with SCExAO/CHARIS222SCExAO stands for the Subaru Coronagraphic Extreme Adaptive Optics Project; CHARIS refers to the Coronagraphic High Angular Resolution Imaging Spectrograph (Jovanovic et al., 2015b; Groff et al., 2016). in the near-infrared (IR), imaging from Keck/NIRC2 in the thermal IR, and astrometric data from the Hipparcos-Gaia Catalogue of Accelerations (HGCA Brandt, 2021). Its spectrum matches that of an L9.5 object: near the L/T transition probing the dissipation of clouds at high altitude in substellar atmospheres (e.g. Saumon & Marley, 2008). Cross-correlating very high spectral resolution data with templates identifies water and carbon monoxide lines in its atmosphere and places some constraints on the companion’s carbon-to-oxygen ratio (Hsu et al., 2024). Joint modeling of the direct imaging data, astrometric data, and archival radial-velocity (RV) constrained the mass to be 46 8 . Updated modeling using Gaia eDR3 astrometry more precisely estimates the companion mass (50 ) and likewise improves orbital constraints (e.g. = 45.2 degrees vs. 39.4 degrees) (Brandt et al., 2021a).
With a near-IR companion-to-star contrast of only 510-5, an angular separation of 07, and clear orbital motion detected over two years, HD 33632 Ab is particularly amenable to follow-up dynamical and atmospheric characterization. As shown in Currie et al. (2021), two epochs of HD 33632 Ab astrometric points from direct imaging separated by 2 years from imaging reduce the companion mass uncertainty by 50% compared to a single epoch of imaging data. Additional epochs may further constrain the companion’s mass and orbital properties. Because HD 33632 Ab is an old, relatively inactive F8V star (Currie et al., 2020a), it is well suited for additional precision RV measurements, complementing astrometric data. The CHARIS data in the discovery paper consist of integral field spectroscopy (IFS) at a very low resolution ( 20). Higher-resolution IFS detections of HD 33632 Ab may enable a more precise characterization of the companion’s atmospheric properties like gravity and carbon chemistry (e.g. Barman et al., 2011).
In this study, we present the first comprehensive atmospheric and dynamical follow-up characterization of HD 33632 Ab, combining new direct imaging and RV data with HGCA astrometry to constrain the companion’s atmosphere, orbit, and mass. Direct imaging data from SCExAO/CHARIS at and band at higher resolution bet: these data and Keck/NIRC2 imaging nearly triple HD 33632 Ab’s astrometric baseline. Similarly, new data from the SOPHIE spectrograph double the RV time baseline compared to data presented in Currie et al. (2020a).
2 Data
2.1 High-Contrast Imaging Observations and Data Reduction
2.1.1 SCExAO/CHARIS Data
We obtained follow-up observations of HD 33632 A with SCExAO/CHARIS in the and band “high resolution” modes (R 60–70) on 18 October 2021 (Table 1). Conditions were photometric and with moderate windspeeds ( 15 mph), resulting in average seeing ( 05–07) for Maunakea. We targeted HD 33632 for a total of 1 hour clock time ranging between 1.2 and 2.3 hours after transit, alternating between the and filters.
All CHARIS data consisted of 60.48 exposures obtained in angular differential imaging(ADI)/pupil tracking mode, allowing the sky to rotate on the detector with time (Marois et al., 2006). The total integration time and parallactic angle motion for and were 27.2 minutes, 19.28o and 25.2 minutes, 14.87o, respectively. To provide astrometric and spectrophotometric calibration, we modulated the SCExAO deformable mirror with an amplitude of 25 nm to generate satellite spots (Jovanovic et al., 2015a). For all data, we inserted a Lyot coronagraph into the focal plane with a 0139 occulting spot diameter, suppressing the stellar halo.
To extract data cubes from the raw CHARIS frames, we used the pipeline from Brandt et al. (2017). Subsequent basic processing followed standard steps utilized for reducing low-resolution/broadband data using the CHARIS Data Processing Pipeline (Currie et al., 2020b). For the K-band data, we used blank sky exposures to subtract thermal background emission. For both passbands, we then registered each cube to a common center and spectrophotometrically calibrated the data assuming a Kurucz atmosphere model appropropriate for an F8V star equal to HD 33632 A’s brightness ( = 5.19 0.02, 5.17 0.02).
Inspection of the extracted CHARIS cubes shows that, like the low-resolution data presented in Currie et al. (2020a), HD 33632 Ab is faintly visible in many individual channels even without point-spread function (PSF) subtraction, especially in band. Combined with the modest parallactic angle rotation enabled by our observations, we therefore tuned our point-spread function (PSF) subtraction approach towards conservative settings resulting in minimal spectrophotometric and astrometric biasing, not to maximize signal-to-noise. We used the Adaptive, Locally-Optimized Combination of Images (A-LOCI) approach (Currie et al., 2012, 2015) employing a pixel mask over the subtraction zone (Marois et al., 2010; Currie et al., 2012) and constructing a reference PSF by minimizing the residuals within the optimization zone (see Lafrenière et al., 2007). This approach results in less signal loss and – when compared with forward-modeling without masking e.g. as with KLIP – significantly reduces the possibility that the companion’s bright signal complicates forward-modeling techniques needed to extract accurate spectrophotometry (Currie et al., 2018). Our reductions used a singular value decomposition (SVD) cutoff of 10-6 and a rotational gap of = 0.75 PSF footprints.
2.1.2 Keck/NIRC2 Data
We also obtained follow-up Keck II imaging of HD 33632 with the NIRC2 data in the broadband filter ( = 3.78 ) using Keck’s facility AO system. Conditions were exceptional for thermal IR AO imaging, with seeing reaching as low as 020, minimal winds, and less than 5% humidity. We used a Lyot coronagraph with a 02 diameter pupil mask and also observed in ADI mode for a total of 25 minutes of integration time and 27o of parallactic angle motion. For photometric calibration, we observed the bright star HD 32537 just prior to our HD 33632 sequences.
We reduced the NIRC2 data using the ADI-based pipeline from Currie et al. (2011). Key steps included sky subtraction, registering each image to a common center, photometric calibration, spatial filtering, PSF subtraction (with A-LOCI), and forward-modeling. Our algorithm settings are marginally more aggressive – identical rotation gap but no pixel masking – as HD 33632 Ab was less clearly visible in raw data than with CHARIS.
| UT Date | Instrument | Filter | [E,N] |
| Previous Data | |||
| 20181018 | SCExAO/CHARIS | ||
| 20181101 | Keck/NIRC2 | ||
| 20200831-20200901 | SCExAO/CHARIS | ||
| New Data | |||
| 20211018 | SCExAO/CHARIS | ||
| 20240223 | Keck/NIRC2 |
2.1.3 Detections
Figure 1 shows the detection of HD 33632 Ab in , , and bands located at 3 o’clock from the star. The estimated signal-to-noise ratios (SNR) for the collapsed cubes in CHARIS and band and NIRC2 filter are 27.6, 29.16, and 29.88 respectively: slightly smaller than the highest-quality broadband CHARIS dataset from Currie et al. (2020a). The first and last channels for each CHARIS data set are contaminated by telluric emission and thus generally have low throughput and low SNR. Aside from these, typical detection significances are SNR 10-25 per channel.
2.2 High-Contrast Imaging Spectral Extraction and Astrometry
To correct for astrometric and spectrophotometric biasing of HD 33632 Ab due to processing, we employed forward-modeling as described in Currie et al. (2018), treating the companion signal as a small perturbation on the LOCI coefficients used to construct the reference PSF (see also Pueyo, 2016). Forward-modeling reveals low signal loss (12-20% per channel) and negligible astrometric biasing. Figure 2 compares the CHARIS and band spectra with the broadband, lower-resolution spectrum from Currie et al. (2020a); Appendix A lists the spectra’s the flux density, uncertainty in the flux density, and SNR in each channel. The spectra agree to within errors at all channels except at 1.8 region affected by telluric absorption.
Figure 3 displays the spectral covariance matrices (Greco & Brandt, 2016) for H (top panel) and K band (bottom panel). The off-diagonal elements represent the spectral and spatial correlation of residuals, identifying spectrally correlated noise. For band, residuals in a given channel are typically correlated with those 2 channels away at a level of 0.5. Covariances are weaker in K band.
We derive and band photometry by convolving the CHARIS and -band spectra with the Mauna Kea Observatories filter transmission (e.g. Currie et al., 2018). HD 33632 Ab’s and band apparent magnitudes are 16.08 0.06 and 15.36 0.06, respectively, and its - color is then 0.72 0.08. For NIRC2 photometry, we used the bright star HD 32537 for photometric calibration, yielding m = 13.64 0.06. These values agree with ones derived from lower-resolution broadband spectra from Currie et al. (2020a) to within 1- for each measurement.
Table 2 lists our astrometry and previous measurements from Currie et al. (2020a). We adopt the astrometric calibration from Currie et al. (2022) appropriate for CHARIS observations obtained through 2021. HD 33632 Ab’s position with CHARIS obtained from averaging the and band astrometry is estimated to be at [E,N] = [0728,0034] [0005,0005]. The astrometric errors consider algorithm biasing (negligible), the intrinsic SNR of the detection, centroiding uncertainty (set to 0.25 pixels or 0004), and differences in the and band measurements. The NIRC2 astrometry was measured to be [E,N] = [0678,0068] [0005,0005]. Compared to its 2020 position, HD 33632 Ab has shifted closer to its host star by a displacement of ([E,N]) [0012,0061] in the CHARIS data and ([E,N]) [0062,0163] in the NIRC2 data.
2.3 Radial-Velocity Data From the SOPHIE Archive
To our high-contrast imaging data, we add precision radial-velocity (RV) data from the SOPHIE instrument (Bouchy & Sophie Team, 2006; Bouchy et al., 2013; Perruchot et al., 2008). The data were taken between 2006 October 14 and 2018 December 6 and extend the RV coverage from Currie et al. (2020a), which was drawn entirely from the Lick Observatory archive data using the Hamilton spectrograph (Fischer et al., 2014). The spectra were taken in high-resolution mode using the Thorium-Argon lamp for wavelength determination. The RV drift caused by a spectrograph instability was monitored by simultaneously observing Thorium-Argon lamp or Fabry-Pérot étalon. Exposure times ranged between 120 and 1200 .
We downloaded the wavelength-calibrated 1D spectra of HD 33632A from the SOPHIE archive333http://atlas.obs-hp.fr/sophie/. To measure RVs of HD 33632 A the spectrum data were analyzed using the SpEctrum Radial Velocity AnaLyser (SERVAL) pipeline (Zechmeister et al., 2018), which can yield a typical RV precision of 1 . When running SERVAL, we ignored spectrum orders in high dispersion spectrscopy if their SNRs are smaller than 10 and set the absolute velocity of HD 33632A to be 1.705 k m-1 according to Soubiran et al. (2018).
Measurements of relative RVs still suffer from imperfect calibrations, varying nightly zero point (NZP) in the RV measurements. The NZP variation of SOPHIE has been monitored by measuring RVs of servarl RV-stable stars in each observing night(Courcol et al., 2015; Hara et al., 2020). Recently, Grouffal et al. (2024) developed a new method to model the NZP drifts by applying Gaussian process (GP) regressions to the RV measurements of the stable stars. The NZP corrections from the GP modeling and the observing epoch corresponding to the NZPs were tablulated in Grouffal et al. (2024), which we adopt in this study. If an observation of HD 33632 A was conducted within one day from an epoch in the NZP table, we adopted the NZP correction of that epoch. Also, the time-nearest NZP correction was adopted if multiple NZP corrections were listed from this selection. We propagated the uncertainties of NZP corrections, which are also tabulated in Grouffal et al. (2024), to the errors of NZP-corrected RVs of HD 33632 A. In contrast, if this selection provided no available NZP correction for an observing epoch, we did not correct for the NZP in that epoch but instead inflated the RV errors by adding the scatter (1.7 m s-1) of all the NZP values in quadrature. Appendix B lists our extracted RVs and uncertainties. The spectra have a mean RV uncertainty of 1.8 . Between October 2014 (JD = 2456939.5) and December 2018 (JD = 2458459.5), the SOPHIE data suggests a long-term RV trend, consistent with a possible RV detection of HD 33632 Ab.
3 Atmospheric Analysis
3.1 Empirical Comparisons
We compared HD 33632 Ab’s H and K band spectra to those from MLT dwarfs in the Montreal Spectral Library (Gagné et al., 2015; Robert et al., 2016; Delorme et al., 2012; Naud et al., 2014). We compare each library spectrum to HD 33632 Ab’s by computing the value as the sum of values in and band, separately:
where we choose the scaling coefficient that minimizes . and are the flux of the template and target spectra, respectively, in the H band wavelength bin or K band wavelength bin. The H and K band spectral covariances are and .
Fig 4 shows that our comparisons reach a minimum for the L9.5 spectral type, where the best-fitting object is the field L9.5 dwarf SIMPJ0956-1447 (Robert et al., 2016), which was also the best fit to the CHARIS low-resolution spectrum from Currie et al. (2020a). The top panel of Figure 5 demonstrates how earlier and later spectral types provide a poorer match to HD 33632 Ab. Consistent with results from Currie et al. (2020a) for low-resolution spectra, HD 33632 Ab’s and band spectra best represent an L9.5 dwarf, where spectral types between L8.5 and L9.5 are favored444Figure 4 shows one L4.5 dwarf also with a low value of 1.3: SIMPJ1122+0343. However, this object’s spectrum’s spectrum is particularly noisy compared to most others in the Montreal Spectral Library. Other L4–L5 dwarfs provide a poor match to HD 33632 Ab’s spectrum..
The bottom panel of Figure 5 compares the HD 33632 Ab and band spectra to this object, the low-resolution HD 33632 Ab CHARIS spectrum from Currie et al. (2020a), and and band spectra for HR 8799 d from Zurlo et al. (2016) and Greenbaum et al. (2018) scaled to HD 33632 Ab. SIMPJ0956-1447 is an excellent match to the HD 33632 Ab and band spectra ( 1.1) and likewise the band portion of the low-resolution spectrum. In contrast, the HR 8799 d spectrum is flatter and redder, fainter at and band but significantly brighter at 2.2 sensitive to methane absorption.
3.2 Atmospheric Modeling
| Model Grid | Metallicity | carbon chemistry | Best Fit Parameters () | ||
| Full Grids | |||||
| BT-Settl/AGS2009 | 400-2500 | 3.5-5.5 | solar | equilibrium chemistry | = 1300 , log(g) = 4.0, = 0.84 (2.973) |
| Exo-REM/S=0.003 | 400-2000 | 3.0-5.0 | [M/H] = -1 to 1 | non-equilbrium/ | = 1250 , log(g) = 5.0, = 0.97 |
| [C/O] = 0.1–0.8 | [M/H] = 0, [C/O] = 0.55 (1.575) | ||||
| Selected Models | |||||
| Lacy/Burrows/AE100 | 1000-1500 | 3.5-5.0 | solar | equilbrium chemistry | = 1400 , log(g) = 4.5, = 0.74 (2.511) |
| Lacy/Burrows/AEE100 | 1000-1500 | 3.5-5.0 | solar | equilbrium chemistry | = 1500 , log(g) = 4.5, = 0.64 (2.386) |
| Lacy/Burrows/AEE100 | 1000-1500 | 3.5-5.0 | solar | non-equilibrium/0.1CH4 | = 1300 , log(g) = 4.5, = 0.8 (2.198) |
| Lacy/Burrows/E60 | 1300-1600 | 4.0-5.0 | solar | non-equilibrium/0.1CH4 | = 1400 , log(g) = 4.5, = 0.70 (2.634) |
To further characterize HD 33632 Ab’s atmosphere, we fit our CHARIS spectra and NIRC2 photometry to atmosphere models varying temperature, gravity, metallicity, and chemistry. Large grids of atmosphere models identify the parameter space best fitting HD 33632 Ab data; selected atmosphere models from other sources explore the robustness of these conclusions to different modeling formalisms. Table 3 summarizes the parameter space covered by each of these model sources, the parameter space they cover, and properties of the best-fit models from each source.
3.2.1 Modeling Approach
The BT-Settl and Exo-REM Model Grids
First, we consider two grids of models: the widely-used BT-Settl models555These models were downloaded from the Theoretical Spectra Web Server: http://svo2.cab.inta-csic.es/theory/newov2/. and the recently-available Exo-REM models (Allard et al., 2012; Charnay et al., 2018). The BT-Settl atmosphere models adopt abundances from Asplund et al. (2009) abundances. Abundances sources for the Exo-REM models include the ExoMol database (Tennyson & Yurchenko, 2012; Yurchenko & Tennyson, 2014) and others described in Baudino et al. (2015, 2017) and Charnay et al. (2018); the Exo-REM models include non-equilibrium carbon chemistry from Charnay et al. (2018), while the BT-Settl models assume equilibrium chemistry.
Both model grids include a self-consistent formalism for clouds and dust entrained in clouds. The Exo-REM grid corresponds to a case with simple microphysics (iron and silicate clouds) with supersaturation parameter S=0.003, which nearly matches the near-IR color-magnitude diagram positions of field L/T transition at high gravity (log(g) = 5) and matches the reddened/underluminous L/T sequence of young, planet-mass companions at low gravity (log(g) = 3.5-4.5) (see Figure 16 in Charnay et al., 2018).
Lacy/Burrows Cloudy Models
Second, we consider selected models from B. Lacy and A. Burrows used in Lacy & Burrows (2020), as updated in in Currie et al. (2023b) to fit HIP 99770 b data. These models cover a more limited range in temperature and gravity than the BT-Settl or Exo-REM grids. But in contrast to these grids, the Lacy/Burrows models parameterize and vary the cloud shape function (i.e. the cloud thicknesses at a given temperature and pressure) and carbon chemistry. Sources for the models’ molecular abundances and pressure-dependent opacities are listed in the Supplementary Material from Currie et al. (2023b)
For our fits, we trim the first and last CHARIS channels for each spectrum to remove those affected by telluric absorption. Following Currie et al. (2023b), we use the statistic to assess the fit quality to the model:
where the vectors and are the difference between measured () and predicted () CHARIS H and K band spectral points () and and are their covariances. The vectors , , and are measured photometry, model predicted photometry, and photometric uncertainty for the measurement.The scaling factor is a free parameter we vary to minimize and is equal to the ratio of the object radius to its distance squared ( (R/D)2 ). We assume a distance of 26.39 (Gaia Collaboration et al., 2023).
| () | Metallicity | C/O | ) | |
| 1250 | 5.0 | 1.0 | 0.55 | 1.575 |
| 1050 | 4.5 | 1.0 | 0.55 | 1.727 |
| 1100 | 5.0 | 1.0 | 0.30 | 1.784 |
3.2.2 Results
Figures 6 compares our HD 33632 Ab data to the best-fitting models from the BT-Settl and Exo-REM grids. The BT-Settl and Exo-REM best-fit models favor a similar temperature (1250–1300 ). The Exo-REM models more accurately reproduce the flattened band shape, yielding a significantly lower minimum reduced of = 1.575 (vs. = 2.973). The best-fit Exo-REM model favors a high surface gravity of log(g) = 5, solar metallicity, and carbon-to-oxygen ratio also near solar ([C/O] = 0.55). The best-fit radius is 0.97 , within 10% of values predicted for the radii of mature brown dwarfs (e.g. Baraffe et al., 2003). most other Exo-REM models with 2 likewise have log(g) = 5.
Table 4 lists the models that match the data to within the 3- confidence limit; Figure 7 shows contours of 2 and 5 for the temperature and gravity for the Exo-REM models at solar metallicity and three different C/O ratios (0.3, 0.55, and 0.7) and the BT-Settl models. The best-fitting Exo-REM models have temperatures of 1100–1350 and high gravities (log(g) 5) or slightly cooler temperatures (1000-1100 ) and lower gravities (log(g) 4–4.5). Low gravity objects with temperatures of 1000-1100 are more consistent with very young ( 100-200 Myr old) planet-mass objects (e.g. Currie et al., 2011; Barman et al., 2015; Currie et al., 2018), while HD 33632 A is at least 1 Gyr old (e.g. Brandt et al., 2021a). The BT-Settl models generally fit the data poorly.
Figure 8 explores the dependency of the fit on carbon chemistry using the Lacy/Burrows models. The best-fitting model has a temperature and gravity comparable to that found from the BT-Settl and Exo-REM grids and is depleted in methane. Equilibrium chemistry models with the same temperature and gravity predict much more methane absorption at band than seen. While equilibrium chemistry models with higher temperatures can much better reproduce the HD 33632 Ab spectrum, their implied radii are significantly smaller, in conflict with predicted radii from cooling models.
4 Dynamical Analysis
We use the open-source, Python-based Markov Chain Monte Carlo (MCMC) package orvara (Brandt et al., 2021b) to fit the mass and orbit of HD33632 Ab. To accurately measure masses and retrieve orbital parameters, orvara integrates a combination of radial velocities (Rvs), relative astrometry and absolute astrometry data from the Hipparcos and Gaia Catalog of Accelerations (HGCA Brandt, 2021). Our new data more than doubles the astrometric baseline for HD 33632 Ab and adds new RV data to better constrain HD 33632 Ab’s properties.
Following previous work (Currie et al., 2020a; Brandt et al., 2021a), we adopt a prior of and absolute astrometry from Hipparcos and Gaia as reported in the HGCA using Gaia eDR3 measurements. We include relative astrometry from SCExAO/CHARIS and NIRC2 as listed in Table 2. We perform two separate fits to the data: one without the inclusion of new RV measurements from SOPHIE and the recently-published companion RV measurement from Hsu et al. (2024) (Simulation 1) and one with these data (Simulation 2)666Hsu et al. (2024) derived the absolute RV of HD 33632A b to be 8 3 km s-1 with a high-resolution spectrum of HD 33632A b. Because orvara employs a companion’s relative RV (which is taken to be a difference between a companion’s absolute RV and a host star’s absolute RV), we calculated the relative RV based on Hsu et al. (2024) for HD 33632A b and Gaia Collaboration et al. (2018) for HD 33632A. We found that including/excluding Hsu et al. (2024) yielded a negligible difference in the companion’s posterior distributions but incorporate this measurement in our simulations for completness. . We compare posterior distributions of the two simulations to assess the sensitivity of mass and orbital parameters derived from astrometry alone. Because it only has a very marginal influence on the posterior distributions for HD 33632 Ab (Currie et al., 2020a), for simplicity we do not model astrometric data from the distant HD 33632 B binary. In both cases, we fit for only one companion (HD 33632 Ab). For both simulations, we treated the first 500 steps as burn-in and thinned the chains by a factor of 50.
For Simulation 1, we run orvara with 25 temperatures, 100 walkers for each temperature and 400,000 steps per walkers. The MCMC converged with a typical autocorrelation timescale of 45 steps per parameter. The fitted parameters along with the priors are detailed in Table A3. The posterior distributions of the orbital parameters and the HD 33632 Ab’s mass are displayed in Figure 9.
For Simulation 2 adding SOPHIE’s RVs data, we run orvara with a total of 25 temperatures, with 100 walkers for each temperature and 500,000 steps per walkers. The MCMC converged with a typical autocorrelation timescale of 45 steps per parameter. We display in Table A4 the fitted parameters and the priors of HD33632 Ab. Figure 10 displays posterior distributions of orbital parameters and the BD’s mass for this simulation.
Comparing the two simulations demonstrates the sensitivity of companion properties derived predominantly from astrometric data vs. those that include more extensive RV coverage. Simulation 1’s results in Figure 9 largely agree with previous analyses albeit with slightly narrower posterior distributions but including some parameters (e.g. companion mass) with multiple peaks. We estimate a new semimajor axis, inclination, eccentricity and companion mass for HD 33632 Ab of au, degrees, , and , respectively. The mass ratio is .
Compared to previous results from Brandt et al. (2021a), the posterior distributions have comparable peaks and slightly narrower 68% confidence intervals, indicating that additional direct imaging astrometry are not substantially improving the precision with which we extract parameters.
Adding the SOPHIE RV data signficantly improves the precision of all extracted dynamical parameters. With these data added, HD33632 Ab’s posterior distributions are consistently about half as wide. The posterior distributions are single peaked for all parameters displayed in Figure 10. The semi-major axis is now au, the inclination is degrees, and the eccentricity is . HD 33632 Ab’s mass and mass ratio are now constrained to within about 5% and 10% precision, respectively: and .
5 Discussion
Analysis of the SCExAO/CHARIS and band spectra presented in this work further clarify HD 33632 Ab’s atmospheric properties beyond the empirical comparisons to lower resolution CHARIS spectra presented in Currie et al. (2020a). HD 33632 Ab’s spectra are well fit by field brown dwarfs with spectral types of L8.5–L9.5. Atmospheric model comparisons find a best-fit temperature of 1250 and favor a high gravity (log(g) = 5.0). The companion’s atmosphere is consistent with having a solar C/O ratio. HD 33632 Ab has a cloudy atmosphere. The best-fitting models (i.e. from the Exo-REM grid) incorporate disequilibrium carbon chemistry; comparisons with the Lacy & Burrows grids show that an atmospheric model with equilibrium carbon chemistry provides poorer fits to the data, especially at band.
Our dynamical analysis illustrates the key role that complementary RV data play even for companion search programs focused on direct imaging and astrometric data. In contrast with earlier HD 33632 results comparing modeling results with single-epoch imaging data vs. two epochs (Currie et al., 2021), we found little improvement in mass and orbit constraints on HD 33632 Ab from additional positional measurements from imaging777It is possible that this result is specific to only a subset of brown dwarfs and planets detected from imaging and astrometry. Analysis of other substellar companions detected with imaging and astrometry yields progressively tighter constraints on mass and orbital properties as additional imaging data are accumulated (T.C., unpublished). . However, when new RV data were incorporating into our modeling, the width of the posterior distributions for HD 33632 Ab’s mass and orbital properties generally shrunk by about a factor of 1.5 to 2 (e.g. a mass of ). RV data offer a fuller representation of companion’s dynamical influence on its host star when combined with astrometry. They should be obtained when possible for other accelerating stars that are targets of planet and brown dwarf searches, especially Sun-like stars amenable to precision RV measurements (e.g. HIP 21152; Kuzuhara et al., 2022; Franson et al., 2023).
Our study provides a complementary probe of HD 33632 Ab’s atmosphere using different data sources than Hsu et al. (2024). They focused on high-dispersion coronagraphic spectroscopy to detect water and carbon monoxide at peak SNRs of 4.8 and 2.7 using the cloudless Sonora Bobcat atmosphere models as molecular templates (Marley et al., 2021) and SNRs of 7.8 and 6.1 using the BT-Settl models and the their baseline forward retrieval model using the petitRadTrans code (Mollière et al., 2019), which both incorporate clouds. Like Hsu et al. (2024), we find evidence for disequilibrium chemistry in HD 33632 Ab’s atmosphere. Temperatures derived from their high-resolution data alone (1550–2500 ) imply spectral types of M6–L4 and are easily ruled out by both our empirical comparisons and atmospheric modeling analysis. However, their retrieval posteriors for temperature incorporating photometry and spectra from Currie et al. (2020a) are in better agreement with our results. Orbital parameters from this work are generally consistent with previous studies (Currie et al., 2020a; Brandt et al., 2021a). Our dynamical mass measurements without RVs are in slight tension with theirs (37 ) and our masses derived with the new SOPHIE RV data rule out their results at the 1.5 level. On the other hand, the usage of Gaia eDR3 data resolves the discrepancy between our parallax posterior distribution and theirs.
HD 33632 Ab is a prime target for atmospheric characterization at passbands complementary to those already probed by ground-based telescopes, especially from data obtained with NASA missions. At 3–5 , HD 33632 Ab has a contrast of 8.6 magnitudes, well within reach for detections with both NIRCam and NIRSPEC (Girard et al., 2022; Ruffio et al., 2023); HD 33632 Ab may also be detectable at 10 with MIRI coronagraphy (Boccaletti et al., 2024). These JWST modes have identified multiple molecules in the atmospheres of young exoplanets or planet-mass companions at the L/T transition (e.g. water, methane, carbon monoxide, carbon dioxide, sodium, and potassium), have constrained the presence of clouds and disquilibrium chemistry, and have more precisely determined bolometric luminosities and radii (e.g. Miles et al., 2023; Boccaletti et al., 2024; Franson et al., 2024). Similar JWST observations for HD 33632 Ab would complement these studies by exploring atmospheric properties for an older field L/T transition object with a dynamical mass measurment.
To predict the detectability of HD 33632 Ab for the Nancy Grace Roman Telescope’s Coronagraphic Instrument (CGI), we computed the companion’s optical contrast for a range of Lacy & Burrows atmosphere models centered on the best-fit model (see also Currie et al., 2023b). As shown in Figure 11 (top panel), deep sodium absorption predicted at 600 nm like makes a HD 33632 Ab detection beyond the capability of CGI in the 575 nm passband to be used for CGI’s technology demonstration. For the best-fit Lacy & Burrows mode, we predict a contrast of 210-10, while a much lower surface gravity object indicative of a young planet mass companion has weaker sodium absorption and thus may be detectable. However, at the 730 nm passband planned for long-slit spectroscopy and 825 nm passband available for imaging, HD 33632 Ab’s predicted contrast is 10-7 and 10-6, respectively. The predicted location for HD 33632 early in the Roman mission (e.g. January 2028) lies beyond the 045 outer working angle for suitable tech demo targets but at the edge of the dark hole region for spectroscopic mode ( 055-06) and well within the dark hole region for 825 nm data (bottom panel). Optical long-slit spectra for HD 33632 Ab cover potassium absorption, a feature whose depth and shape is expected to be sensitive to metallicity. The full optical SED of HD 33632 Ab may also probe other atmospheric properties like clouds (Lacy & Burrows, 2020).
Finally, upcoming instruments behind ground-based extreme AO systems may similarly provide additional information on HD 33632 Ab’s atmospheric properties, especially those operating at higher spectral resolution than CHARIS. Hsu et al. (2024) demonstrated the ability of high-resolution spectroscopy to resolve HD 33632 Ab’s spectral lines and constrain rotation. Medium-resolution spectroscopy in combination with molecular mapping can yield higher SNR detections of individual molecules from applying a cross-correlation function to theoretical templates and companion spectra (e.g. Hoeijmakers et al. (2018)) without relying on a precise fiber alignment to HD 33632 Ab’s position.
The Exo-NINJA medium-resolution integral field spectrograph at Subaru (El Morsy et al., 2024) will provide this capability from the red optical through K band and will sit behind AO3K: the extreme AO upgrade to Subaru’s facility AO system (Lozi et al., 2022). Future observations with extremely-large telescopes (e.g. Thirty Meter Telescope) will provide even higher quality spectra. Robust atmospheric characterization of HD 33632 Ab from the optical through 10 will cement it as a key object for understanding the atmospheres of substellar objects of a given mass and a reference point for understanding exoplanetary atmospheres.
The development of SCExAO was supported by JSPS (Grant-in-Aid for Research #23340051, #26220704 & #23103002), Astrobiology Center of NINS, Japan, the Mt Cuba Foundation, and the director’s contingency fund at Subaru Telescope. CHARIS was developed under the support by the Grant-in-Aid for Scientific Research on Innovative Areas #2302. Some of the data presented herein were obtained at the W. M. Keck Observatory, which is operated as a scientific partnership among the California Institute of Technology, the University of California and the National Aeronautics and Space Administration. The Observatory was made possible by the generous financial support of the W. M. Keck Foundation.
T.C. is supported by NASA-Keck Strategic Mission Support grant 80NSSC24K0943. M.K. is supported by JSPS KAKENHI grant No. 24K07108. M.T. is supported by JSPS KAKENHI grant No. 24H00242.
Appendix A HD 33632 Ab CHARIS Spectra
| Wavelength (m) | (mJy) | (mJy) | SNR |
| 1.48 | 0.236 | 0.049 | 5.0 |
| 1.49 | 0.214 | 0.025 | 9.2 |
| 1.51 | 0.219 | 0.021 | 10.7 |
| 1.52 | 0.301 | 0.019 | 16.5 |
| 1.54 | 0.292 | 0.022 | 13.6 |
| 1.55 | 0.377 | 0.020 | 20.6 |
| 1.57 | 0.369 | 0.020 | 19.5 |
| 1.59 | 0.447 | 0.020 | 24.2 |
| 1.6 | 0.431 | 0.019 | 24.0 |
| 1.62 | 0.448 | 0.020 | 24.1 |
| 1.63 | 0.428 | 0.021 | 21.5 |
| 1.65 | 0.483 | 0.022 | 23.3 |
| 1.67 | 0.444 | 0.023 | 20.5 |
| 1.69 | 0.496 | 0.019 | 28.3 |
| 1.7 | 0.445 | 0.029 | 15.8 |
| 1.72 | 0.442 | 0.019 | 24.5 |
| 1.74 | 0.374 | 0.021 | 18.5 |
| 1.76 | 0.339 | 0.017 | 21.8 |
| 1.77 | 0.309 | 0.035 | 11.3 |
| 1.79 | 0.123 | 0.160 | 0.8 |
| 2.02 | 0.297 | 0.038 | 9.3 |
| 2.04 | 0.414 | 0.017 | 27.7 |
| 2.06 | 0.455 | 0.023 | 23.2 |
| 2.08 | 0.492 | 0.025 | 22.9 |
| 2.1 | 0.519 | 0.023 | 25.0 |
| 2.12 | 0.534 | 0.022 | 29.5 |
| 2.14 | 0.568 | 0.029 | 21.6 |
| 2.16 | 0.52 | 0.025 | 23.5 |
| 2.18 | 0.526 | 0.032 | 18.0 |
| 2.21 | 0.465 | 0.030 | 17.0 |
| 2.23 | 0.489 | 0.038 | 13.7 |
| 2.25 | 0.461 | 0.040 | 12.2 |
| 2.27 | 0.428 | 0.049 | 9.2 |
| 2.3 | 0.404 | 0.084 | 5.0 |
| 2.32 | 0.278 | 0.056 | 5.1 |
| 2.34 | 0.423 | 0.076 | 5.8 |
| 2.37 | 0.45 | 0.073 | 6.9 |
Note. — Throughput of HD 33632 Ab spectrum in H and K bands extracted from 18 October 2021 and reduced by A-LOCI.
Appendix B New SOPHIE Radial-Velocity Measurements
| MBJD | RV (m/s) | RVerr (m/s) |
| 54023.135 | -1.809 | 2.535 |
| 54023.145 | -4.324 | 2.421 |
| 54024.098 | 6.313 | 2.351 |
| 54026.138 | 4.961 | 2.555 |
| 54086.061 | 9.429 | 3.110 |
| 54089.965 | 9.092 | 2.289 |
| 54096.954 | 6.304 | 2.108 |
| 54186.829 | 30.722 | 3.003 |
| 54375.167 | 13.564 | 2.263 |
| 54496.821 | 33.295 | 2.229 |
| 54500.781 | 43.278 | 2.282 |
| 54501.797 | 36.277 | 2.602 |
| 54502.813 | 47.435 | 2.469 |
| 54504.802 | 17.764 | 2.527 |
| 54505.812 | 19.926 | 2.284 |
| 54506.821 | 27.577 | 2.221 |
| 54718.146 | 29.359 | 2.531 |
| 54724.106 | 36.429 | 2.807 |
| 54725.156 | 38.092 | 2.438 |
| 54726.167 | 22.705 | 2.158 |
| 54729.153 | 26.497 | 2.074 |
| 54733.165 | 17.943 | 2.398 |
| 54734.136 | 26.257 | 2.430 |
| 56939.063 | 2.780 | 1.373 |
| 57063.834 | -2.120 | 1.475 |
| 57080.880 | 7.210 | 1.540 |
| 57289.088 | -11.610 | 1.584 |
| 57342.111 | -17.439 | 1.510 |
| 57378.067 | -18.150 | 1.483 |
| 57416.955 | -63.463 | 13.328 |
| 57468.849 | -25.023 | 1.303 |
| 57637.133 | -25.530 | 1.422 |
| 57664.165 | -24.284 | 1.860 |
| 57729.130 | -30.735 | 0.953 |
| 58051.112 | -33.599 | 1.526 |
| 58073.015 | -27.343 | 1.373 |
| 58130.871 | -30.509 | 1.168 |
| 58214.805 | -24.676 | 1.535 |
| 58351.140 | -43.808 | 2.100 |
| 58457.026 | -43.470 | 1.364 |
| 58457.944 | -28.540 | 1.417 |
| 58459.023 | -44.885 | 1.364 |
Appendix C Tables of orvara Modeling Results
| Parameter | 16/50/84% quantiles | Prior |
| Fitted Parameters | ||
| RV jitter (m/s) | log-flat | |
| () | Gaussian, | |
| () | (log-flat) | |
| Semimajor axis (au) | (log-flat) | |
| **Orbital parameters listed are of HD 33632 Ab. HD 33632 Aa has shifted by 180∘. | uniform | |
| **Orbital parameters listed are of HD 33632 Ab. HD 33632 Aa has shifted by 180∘. | uniform | |
| Inclination (∘) | (geometric) | |
| PA of the ascending node (∘) | uniform | |
| Mean longitude at 2010.0 (∘) | uniform | |
| Parallax (mas) | Gaussian, 0 | |
| Derived Parameters | ||
| Period (yrs) | ||
| Argument of periastron (∘)**Orbital parameters listed are of HD 33632 Ab. HD 33632 Aa has shifted by 180∘. | ||
| Eccentricity | ||
| Semimajor axis (mas) | ||
| Periastron time (JD) | ||
| Mass ratio | ||
| Parameter | 16/50/84% quantiles | Prior |
| Fitted Parameters | ||
| RV jitter (m/s) | , | log-flat |
| () | Gaussian, | |
| () | (log-flat) | |
| Semimajor axis (AU) | (log-flat) | |
| **Orbital parameters listed are of HD 33632 Ab. HD 33632 Aa has shifted by 180∘. | uniform | |
| **Orbital parameters listed are of HD 33632 Ab. HD 33632 Aa has shifted by 180∘. | uniform | |
| Inclination (∘) | (geometric) | |
| PA of the ascending node (∘) | uniform | |
| Mean longitude at 2010.0 (∘) | uniform | |
| Parallax (mas) | Gaussian, 0 | |
| Derived Parameters | ||
| Period (yrs) | ||
| Argument of periastron (∘)**Orbital parameters listed are of HD 33632 Ab. HD 33632 Aa has shifted by 180∘. | ||
| Eccentricity | ||
| Semimajor axis (mas) | ||
| Periastron time (JD) | ||
| Mass ratio | ||